A water main for a street is being laid using particular kind of pipe that comes in either 18-foot sections or 20-foot sections. The designer has determined that the water main would require 14 fewer sections of 20-foot pipe than if 18-foot sections were used. Find the total length of the water main.
Found 2 solutions by stanbon, edjones:
Answer by stanbon(75887)

A water main for a street is being laid using particular kind of pipe that comes in either 18-foot sections or 20-foot sections. The designer has determined that the water main would require 14 fewer sections of 20-foot pipe than if 18-foot sections were used. Find the total length of the water main.
--------------------
Let the total length of the water main be x feet.
# of 18 ft - # of 20 ft = 14
x/18 - x/20 = 14
(20x-18x)/20*18 = 14
2x = 20*18*14
x = 10*18*14
x = 2520 ft.
==========================
Cheers,
Stan H.

Answer by edjones(8007)

Let x be the number of sections of 18' pipe.
main length=pipe length * number of sections of pipe
20(x-14)=18x
20x-280=18x
20x-18x-280+280=18x-18x+280
2x=280
2x/2=280/2
x=140
18*140=2520' Length of main.
Check:
20(140-14)=20*126=2520
.
Ed